(2x+3y)dx-4(x+1)dy=0
1
Expert's answer
2020-12-11T11:49:04-0500

Let us solve the differential equation "(2x+3y)dx-4(x+1)dy=0". It follows that "x=-1" is a solution. Further, we will assume that "x\ne -1". So, in this case the equation is equivalent to the equation


"\frac{dy}{dx}=\frac{2x+3y}{4(x+1)}".


Let us solve the system "\begin{cases}2x+3y=0\\x+1=0\end{cases}" which is equivalent to "\begin{cases}x=-1\\y=\frac{2}{3}\end{cases}".


Let us introduce a substitution "\begin{cases}x=t-1\\y=z+\frac{2}{3}\end{cases}".


Then we have he differential equation "\frac{dz}{dt}=\frac{2(t-1)+3(z+\frac{2}{3})}{4(t-1+1)}=\frac{2t+3z}{4t}".


Let us introduce a substitution "z=t\cdot u(t)". Then "dz=u\cdot dt+t\cdot du" and thus


"\frac{u\cdot dt+t\cdot du}{dt}=\frac{2t+3tu}{4t}= \frac{2+3u}{4}"


"u+t\frac{du}{dt}= \frac{2+3u}{4}"


"t\frac{du}{dt}= \frac{2+3u}{4}-u=\frac{2-u}{4}"


If "u=2", then "z=2t", "y-\frac{2}{3}=2x+2", and therefore, "y=2x+2+\frac{2}{3}=2x+\frac{8}{3}". Since "(2x+3y)dx-4(x+1)dy=(2x+3(2x+\frac{8}{3}))dx-4(x+1)2dx=(8x+8)dx-8(x+1)dx=0," we conclude that "y=2x+\frac{8}{3}" is a solution of the differential equation.


Let "u\ne 2". Then


"\frac{4du}{2-u} =\frac{dt}{t}"


"\int\frac{4du}{2-u} =\int\frac{dt}{t}"


"-4\ln|2-u|+\ln|C|=\ln|t|"


"\ln|C|=\ln|t|+ \ln(2-u)^4=\ln|t(2-u)^4|"


"t(2-u)^4=C"


"t(2-\frac{z}{t})^4=C"


"(x+1)(2-\frac{y-\frac{2}{3}}{x+1})^4=C"


"(x+1)(\frac{2x-y+\frac{8}{3}}{x+1})^4=C"


"(2x-y+\frac{8}{3})^4=C(x+1)^3"


If "C=0", then we have the solution "y=2x+\frac{8}{3}".


Therefore, we have the following solution of the differentail equation "(2x+3y)dx-4(x+1)dy=0":


"(2x-y+\frac{8}{3})^4=C(x+1)^3" and "x=-1."